OpenStudy (anonymous):

factor: f(x) = x^2-1/3x-1/18

OpenStudy (anonymous):

f(x)=18*(x^2-1/3x-1/18) = 18x^2 -6x -1 = [-(-6) ∓ √(-6)^2 - (4* 18* -1)]/(2*18) = [ 6 ∓ √36 + 72]/36 = [ 6 ∓ √108]/36 = [ 6 ∓ 6√3]/36 = [ 1 ∓ √3]/6 = ( 1 + √3)/6 or ( 1 - √3)/6

OpenStudy (anonymous):

x=negative b, plus or minus the square root of b squared minus four a(c), ALL over two a. so in this case a is one, b is negative one third, and c is negative one eighteenth. this yields one plus or minus (root 3 over six)